• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Charles Hyde-Wright

    Log entry time 11:37:36 on February 1,2002

    Entry number 81469

    keyword=HRS Motion

    Last evening, we moved HRS-L from 14.57 deg to 4A endpoint kinematics
    40.97 deg. After brief endpoint runs we moved HRS-L to
    42.58 deg for production running.
    Both moves were done without visual inspection in the hall.
    Probably between 40.97 and 42.589 we pushed the concrete stop
    on the floor and snagged a cable bundle. This bundle runs
    from HRS photon arm, in front of the front bogies and to the
    RCS electronics hut. The bundle consists of on HV cable, two
    high quality signal cables and aprox 4 Rg-58.
    The HV and 3/4 RG58 cables jammed under the guard plate
    surrounding the bogie on the large angle side of the spectrometer.
    The cables were stretched very tight (wrapping around a beam
    dump support post). The HV set point of 500 V did not trip.
    The HV cable drives a PMT monitoring the gain monitor laser
    on photon arm. The signal cables are cosmic trigger, laser trig?, etc.

    S0 trigger cable was also coiled dangerously close to middle bogies,
    but untangled

    At 10:30 we went in to install post radiator in front of sweep magnet.
    Could not pull cables free of bogie. Cut cables, and pulled cables
    free during and after motion from 42.58 to 40.97 degrees.
    Beam restore 11:30

    Recommended action:
    Do not make large angle changes without visual inspection.